OpenStudy (rissyroo118):

Which ordered pair is a solution to the equation? 2y = −5x – 4 (Points : 1) (−2, 3) (1, −4) (0, 2) (−1, 1)

OpenStudy (rissyroo118):

helppppp

OpenStudy (domebotnos):

2(3) = -5(-2) - 4 6 = 10 - 4 6 = 6 The answer is A. All u do is plug in. Do u understand? :)

OpenStudy (domebotnos):

Medal, plz? :D

OpenStudy (rissyroo118):

thank you :)

OpenStudy (domebotnos):

Np. :)

OpenStudy (rissyroo118):

could you help m with some more math problems @domebotnos

OpenStudy (domebotnos):

If u still have them I could try.